# Dimension reduction via score ratio matching

### v1

@@ 1 Introduction
In this paper, we are interested in two types of low-dimensional structure that appear in many Bayesian inference problems, and more generally in probabilistic modeling. First is the notion that the target distribution can be well approximated as a low-dimensional update of a dominating reference distribution. Second is the notion that the conditioning variables can be replaced with low-dimensional projections or summaries. A recent line of work (Brennan et al., 2020; Zahm et al., 2022; Cui and Zahm, 2021; Baptista et al., 2022; Cui and Tong, 2022) has developed gradient-based methods for identifying and exploiting both types of low-dimensional structure in non-Gaussian settings. These approaches construct specific diagnostic matrices containing averaged gradient or Hessian information of the posterior or joint log-density. Eigendecompositions of these diagnostic matrices reveal informative projections of the conditioning variables and informed projections of the parameters, and also yield bounds on the error of the resulting approximations to the posterior in terms of the rank of these projections. Specifically, the spectra of these matrices indicate to what extent a given problem has the posited low-dimensional structure.
@@
Related work: Many previous papers have highlighted the benefits of gradient-based dimension reduction in Bayesian computation, for methods ranging from MCMC (Cui and Tong, 2022; Cui et al., 2014b; Constantine et al., 2016) to SVGD (Chen and Ghattas, 2020) to normalizing flows (Brennan et al., 2020; Cui et al., 2023; Radev et al., 2020) to ensemble filtering (Le Provost et al., 2022). The present work is concerned with realizing these benefits in gradient-free settings, which include Bayesian inverse problems with complex forward models (Kaipio and Somersalo, 2006; Biegler et al., 2010), goal-oriented inference  (Lieberman and Willcox, 2013; Berger et al., 1999), simulation-based inference (SBI) more broadly (Cranmer et al., 2020), and the purely data-driven setting where only a fixed set of samples from $\pi_{X,Y}$ are given. The second form of dimension reduction we seek (low-dimensional conditioning) is related to SBI-focused efforts of creating summary (ideally sufficient) statistics of the observations to improve the quality and efficiency of inference algorithms (Fearnhead and Prangle, 2012; Joyce and Marjoram, 2008; Nunes and Balding, 2010). So far, most of these methods have not been gradient-based, and do not explicitly take advantage of low-dimensional structure when learning statistics. For instance, Radev et al. (2020) learns an (arbitrary) summary neural network to compress the observation variable before it enters an inference network. Brehmer et al. (2020); Alsing et al. (2018) define data summaries based on locally sufficient statistics, but only around one reference parameter value. More generally, hand-crafted summary statistics are often the norm in SBI (Sisson et al., 2018). These statistics often require expert knowledge to form or use specific selection criteria that are not tied to posterior quality, and hence may lead to issues with “poor quality features,” as seen in Lueckmann et al. (2017). In contrast, the approach proposed here seeks optimal summaries that provide error guarantees on the resulting posterior approximation.
@@ 2.1 Gradient-based dimension reduction
We provide a brief description of two gradient-based dimension reduction methods for which we will develop gradient-free versions. These methods are data-free certified dimension reduction (CDR) [^1] We drop ‘data-free’ in the acronym as we will not consider the dimension reduction problem initially presented in Zahm et al. (2022) for the posterior corresponding to a single realization of the observation. (Cui and Zahm, 2021) and conditional-mutual information based dimension reduction (CMIDR) (Baptista et al., 2022). Both methods decompose the parameter and observation into two subspaces,
@@
Proposition 2.2 (Modified from Section 3.3 of Cui and Zahm (2021)).
@@ 2.2 Approximating score functions
Score matching has recently appeared as a powerful unsupervised learning framework with applications to generative modeling (Song and Ermon, 2019; Song et al., 2021; Ho et al., 2020; De Bortoli et al., 2021) and Bayesian inference (Zhang et al., 2018; Pacchiardi and Dutta, 2022). The core task is to approximate the score function, the gradient of a log-density function, for various downstream tasks. We direct readers to Song and Ermon (2019); Song et al. (2021) for an overview of score matching, especially the derivation of the objective function for learning the score, network training strategies, and its application to generative modeling using Langevin sampling. contributions of the latter sections. In this section, we focus on conditional score-matching, where one approximates the score function of the conditional distribution $\nabla_{x}\log\pi_{X|Y}(x|y)$.
@@
While minimizing this objective is intractable as it requires access to the score function, approaches known as implicit score matching (Hyvärinen and Dayan, 2005) and denoising score matching (Vincent, 2011) introduce reformulated objective functions for $J^{*}(\theta)$ to learn the score based only on samples from a given distribution, such as $\pi_{X,Y}$. We will focus on implicit score matching in this work, although this is not a limitation. Under mild regularity assumptions on the true and approximate score functions (see Assumption 3.1), Hyvärinen and Dayan (2005) showed that the objective $J^{*}$ can be written as
@@ 3 Score ratio matching
$J(w_{\theta})\coloneqq\frac{1}{N}\sum_{j=1}^{N}\frac{1}{2}w_{\theta}(x^{(j)},y^{(j)})^{\top}w_{\theta}(x^{(j)},y^{(j)})+\operatorname{Tr}(\nabla_{x}w_{\theta}(x^{(j)},y^{(j)}))+\nabla_{x}\log\rho(x^{(j)})^{\top}w_{\theta}(x^{(j)},y^{(j)}).$ (8)
@@
See Section A.3 for the proof.
@@ 5 Deflating score-ratio matching
Let $(\lambda_{i},w_{i})$, for $i=1,\dots,d$, be the eigenpairs of a symmetric matrix $A\in\mathbb{R}^{d\times d}$, with $\lambda_{1}>\lambda_{2}\geq\dots\lambda_{d}$ and $\|w_{i}\|_{2}=1$. Define the subunitary matrix $W_{r}=[w_{1}\ \dots\ w_{r}]$ and projector $P=I-W_{r}W_{r}^{\top}$. Then the deflated matrix $\widetilde{A}\coloneqq PAP$ has eigenpairs $(0,w_{i})$ for $i=1,\dots,r$ and $(\lambda_{i},w_{i})$ for $i=r+1,\dots,d$.
@@ 6 Numerical examples
We now present several numerical experiments to show the utility of our methods. Table 1 summarizes all network and training hyperparameter choices for each numerical example. For the problems presented in Sections 6.1 and 6.2, the score ratio is tractable and thus we can construct the true diagnostic matrices $H^{X}_{\operatorname{CDR}}$ and $H^{Y}_{\operatorname{CMI}}$. We thus evaluate the accuracy of our method by comparing the posterior approximation errors for the optimal basis transformations in (4) and (6) with the following error bounds achieved using the learned bases $\widetilde{U}$ and $\widetilde{V}$ [^3] We use a tilde to distinguish learned bases, i.e., eigenvectors of the diagnostic matrices computed using a score (ratio) network, from the eigenvectors of the exact diagnostic matrices.:
@@ 6.2 PDE-constrained inverse problem
Next, we consider an inverse problem where the forward model involves an elliptic partial differential equation (PDE). The inference parameter describes the permeability field on a two-dimensional domain $\mathcal{D}$ and the observations are pointwise measures of the corresponding pressure field. This so-called “Darcy flow” problem is a widely-used test case in the literature on nonlinear Bayesian inverse problems (Stuart, 2010; Iglesias et al., 2014; Cui et al., 2014a). The permeability field $e^{\varrho}$ and the pressure field $u$ are related by the following Poisson equation (Neuman and Yakowitz, 1979; McLaughlin and Townley, 1996; Carrera et al., 2005; Sun, 2013):
@@ 6.3 Flux, a quantity-of-interest inference example
The posterior distribution of interest in this problem is $\pi_{Q|Y}$, which follows from the joint distribution $\pi_{Q,Y}$. We seek to reduce the dimension of the observations $Y$ via Algorithm 1 using $N=10000$ samples. Figure 6(a) shows bounds on posterior approximation error that are estimated using our learned score-ratio network. The fast decay in this bound suggests that the posterior should be well approximated using relatively low-dimensional observation projections, e.g., $s=4$. To validate the utility of the corresponding learned basis $\widetilde{V}_{s}$, we perform inference using a conditional normalizing flow (specifically an unconstrained monotonic neural network (UMNN) as described in Wehenkel and Louppe (2019)) trained from samples of $\pi_{Q,Y}$; see Section B.2 for the implementation details. We build two such flows: one depending on the full-dimensional observations and the other depending on the reduced observations of dimension $s=4$ defined by the basis $\widetilde{V}_{s}$. We compare the performance of the UMNN flows learned with an increasing number of map training samples (independent from the samples used to learn the score ratio approximation) to true posterior predictive samples computed via MCMC. [^4] We perform MCMC on the high-dimensional latent parameters $X$ described in Section 6.2, and then compute log-flux samples $Q$ via (12).
@@ 6.4 Energy price modeling
We use this dataset to estimate the score ratio for $\pi_{X|Y}$ and thus the diagnostic matrix $H^{Y}_{\operatorname{CMI}}$ via Algorithm 1. In Figure 6(c), we observe a sharp drop in the estimated posterior error from observation dimension reduction, suggesting that the leading $s=4$ modes should capture most of the information carried by the full observation about the parameter. Figures 7(b) and 7(c) plot the two leading basis vectors for the observation, and labels the observation components with the highest magnitudes. We see that the first vector primarily captures the temperature predictions at a few key cities. The second vector gives the most weight to the temporal encoding. Figure 9 plots the temperature and cloud coverage contribution of the leading four basis vectors on a map of the United States. We see that many vectors have interpretable structures. For example, the temperature component of $v_{3}$ seems to be a weighted average of the temperature predictions across the weather stations.
(a) Flux: bound on the posterior error
(b) Flux: KS statistic vs. map training set size
(c) Energy price: bound on the posterior error
(d) Energy price: CRPS vs. map training set size
Figure 6: Flux and energy price problems: (a/c) decay of the estimated error bounds that may used used to select the dimension of the reduction observation; (b/d) average KS statistic and CRPS versus the number of training samples, respectively. Reducing the dimension of the observation enables higher-fidelity inference.
@@
(a) Example posterior $1$
(b) Example posterior $2$
(c) Example posterior $3$
(d) Example posterior $4$
(e) Example posterior $5$
(f) Example posterior $6$
@@
As before, we validate the learned reduction vectors by comparing the performance of approximate inference using full- and reduced-dimensional conditional normalizing flows as a function of the number of training samples from $\pi_{X,Y}$. While the ground-truth posterior distributions are not available in this problem, we do have the “true” realized value of the energy price. Thus we report the continuous ranked probability score (CRPS) (Gneiting and Raftery, 2007) for the approximate posterior samples, a predictive metric used to evaluate the performance of probabilistic forecasts, on average over $100$ held-out energy-price samples. Figure 6(d) shows that the reduced-dimensional flows are more predictive on average than the full-dimensional flow. We also note that the CRPS is stable with as few as $100$ training samples for the reduced-dimensional map. Figure 8 shows six histograms of posterior samples generated from full- and reduced-dimensional flows trained with $10000$ samples for different realizations of the observation variable. In two of the examples, we see that the reduced flows may capture underlying multi-modality in the posterior that the full-dimensional flow (i.e., without observation reduction) does not.

### v2

@@ 1 Introduction
In this paper, we are interested in two types of low-dimensional structure that appear in many Bayesian inference problems, and more generally in probabilistic modeling. First is the notion that the target distribution can be well approximated as a low-dimensional update of a dominating reference distribution. Second is the notion that the conditioning variables can be replaced with low-dimensional projections or summaries. A recent line of work (Brennan et al., 2020; Zahm et al., 2022; Cui & Zahm, 2021; Baptista et al., 2022; Cui & Tong, 2022) has developed gradient-based methods for identifying and exploiting both types of low-dimensional structure in non-Gaussian settings. These approaches construct specific diagnostic matrices containing averaged gradient or Hessian information of the posterior or joint log-density. Eigendecompositions of these diagnostic matrices reveal informative projections of the conditioning variables and informed projections of the parameters, and also yield bounds on the error of the resulting approximations to the posterior in terms of the rank of these projections. Specifically, the spectra of these matrices indicate to what extent a given problem has the posited low-dimensional structure.
@@
Related work: Many previous papers have highlighted the benefits of gradient-based dimension reduction in Bayesian computation, for methods ranging from MCMC (Cui & Tong, 2022; Cui et al., 2014; Constantine et al., 2016) to SVGD (Chen & Ghattas, 2020) to normalizing flows (Brennan et al., 2020; Cui et al., 2023; Radev et al., 2020) to ensemble filtering (Le Provost et al., 2022). The present work is concerned with realizing these benefits in gradient-free settings, which include Bayesian inverse problems with complex forward models (Kaipio & Somersalo, 2006; Biegler et al., 2010), goal-oriented inference  (Lieberman & Willcox, 2013; Berger et al., 1999), simulation-based inference (SBI) more broadly (Cranmer et al., 2020), and the purely data-driven setting where only a fixed set of samples from $\pi_{X,Y}$ are given.
The second form of dimension reduction we seek (low-dimensional conditioning) is related to SBI-focused efforts of creating summary (ideally sufficient) statistics of the observations to improve the quality and efficiency of inference algorithms (Fearnhead & Prangle, 2012; Joyce & Marjoram, 2008; Nunes & Balding, 2010). So far, most of these methods have not been gradient-based, and do not explicitly take advantage of low-dimensional structure when learning statistics. For instance, Radev et al. (2020) learns an (arbitrary) summary neural network to compress the observation variable before it enters an inference network. Brehmer et al. (2020); Alsing et al. (2018) define data summaries based on locally sufficient statistics, but only around one reference parameter value. More generally, hand-crafted summary statistics are often the norm in SBI (Sisson et al., 2018). These statistics often require expert knowledge to form or use specific selection criteria that are not tied to posterior quality, and hence may lead to issues with “poor quality features,” as seen in Lueckmann et al. (2017). In contrast, the approach proposed here seeks optimal summaries that provide error guarantees on the resulting posterior approximation.
@@ 2.1 Gradient-based dimension reduction
We provide a brief description of two gradient-based dimension reduction methods for which we will develop gradient-free versions. These methods are data-free certified dimension reduction (CDR) [^1] We drop ‘data-free’ in the acronym as we will not consider the dimension reduction problem initially presented in Zahm et al. (2022) for the posterior corresponding to a single realization of the observation. (Cui & Zahm, 2021) and conditional-mutual information based dimension reduction (CMIDR) (Baptista et al., 2022). Both methods decompose the parameter and observation into two subspaces,
@@
Proposition 2.2 (Modified from Section 3.3 of Cui & Zahm (2021)).
@@ 2.2 Approximating score functions
Score matching has recently appeared as a powerful unsupervised learning framework with applications to generative modeling (Song & Ermon, 2019; Song et al., 2021; Ho et al., 2020; De Bortoli et al., 2021) and Bayesian inference (Zhang et al., 2018; Pacchiardi & Dutta, 2022). The core task is to approximate the score function, the gradient of a log-density function, for various downstream tasks. We direct readers to Song & Ermon (2019); Song et al. (2021) for an overview of score matching, especially the derivation of the objective function for learning the score, network training strategies, and its application to generative modeling using Langevin sampling. In this section, we focus on conditional score-matching, where one approximates the score function of the conditional distribution $\nabla_{x}\log\pi_{X|Y}(x|y)$.
@@
While minimizing this objective is intractable as it requires access to the score function, approaches known as implicit score matching (Hyvärinen & Dayan, 2005) and denoising score matching (Vincent, 2011) introduce reformulated objective functions for $J^{*}(\theta)$ to learn the score based only on samples from a given distribution, such as $\pi_{X,Y}$. We will focus on implicit score matching in this work, although this is not a limitation. Under mild regularity assumptions on the true and approximate score functions (see Assumption 3.1), Hyvärinen & Dayan (2005) showed that the objective $J^{*}$ can be written as
@@ 3 Score ratio matching
$J(w_{\theta})\coloneqq\frac{1}{N}\left[\sum_{j=1}^{N}\frac{1}{2}w_{\theta}(x^{(j)},y^{(j)})^{\top}w_{\theta}(x^{(j)},y^{(j)})+\operatorname{Tr}(\nabla_{x}w_{\theta}(x^{(j)},y^{(j)}))+\nabla_{x}\log\rho(x^{(j)})^{\top}w_{\theta}(x^{(j)},y^{(j)})\right].$ (8)
@@
See Section A.3 for the proof. We note that the approximation error of the score ratio network $\epsilon$ is generally unknown in practice. Theorem 3.3, however, establishes stability of the error in the approximate posterior distribution $\widetilde{\pi}_{X|Y}$, which depends on the subspace $U$, with respect to error in the score ratio used to identify that subspace.
Choice of the KL divergence as the error metric
We note that our methods specifically seek projections of the parameter and observation that minimize bounds on the posterior approximation error measured by the data-averaged KL divergence $\mathbb{E}_{\pi_{Y}}\left[D_{\textrm{KL}}(\pi_{X|Y}||\widetilde{\pi}_{X|Y})\right]$, directly following the results of Propositions 2.2 and 2.3. While the data-averaged KL divergence is a natural and widely used choice, alternative error metrics such as a worst-case KL divergence (over $Y$), Wasserstein distances, or total variation distances may provide additional insights or be better suited for specific inference tasks. Developing methods that minimize these alternative metrics and studying the differences in the posterior approximations could be an interesting direction for future work, particularly in cases where robustness to tail behavior or accuracy of specific posterior expectations are important.
Comparison with other common dimension reduction methods
Here we comment on the differences between the present work and several common dimension reduction methods. The classical method of principal component analysis (PCA) seeks directions that maximize the variance of the data. Nonlinear methods such as t-SNE (Van der Maaten & Hinton, 2008) and UMAP (McInnes et al., 2018) can effectively preserve local neighborhood structures, making them powerful tools for visualization and marginal dimension reduction. However, these methods do not explicitly consider the joint structure between parameters and observations—which is crucial for inference problems, where preserving conditional relationships is essential. We reiterate that our framework seeks projections of the parameter and observation that are informed and informative, respectively, making it particularly suited to inference. In comparison, methods that act marginally (e.g., performing PCA, t-SNE, UMAP etc. on $X$ and $Y$ individually) do not capture the dependence structure between the parameter and observation and thus do not explicitly consider the underlying inference problem.
We note that in the case of linear-Gaussian likelihood models, the projections proposed by Propositions 2.2 and 2.3 are equivalent to those found by canonical correlation analysis (CCA) based on covariance information between $X$ and $Y$; see Proposition 5 in Baptista et al. (2022). Insofar as the gradient-based methods proposed in Baptista et al. (2022) can be viewed as a generalization of CCA to non-Gaussian likelihood models (e.g., by using gradient information of a possibly nonlinear forward model), our methods further generalize to problems where gradients of these likelihood functions are not available.
Choice of reference distribution
We emphasize that the choice of the reference distribution in our method is a degree of freedom, provided that its score function is easily evaluated. Our method identifies directions in parameter space where the posterior differs from the reference distribution on average, and is most effective when these directions are contained within a low-dimensional subspace. The marginal distribution of the parameter (i.e., the prior in Bayesian settings) is thus the natural first choice, when tractable. In cases where the prior can be easily mapped to a standard Gaussian, we recommend reparameterizing the problem accordingly for computational convenience (as done in Sections 6.1 and 6.2). However, when such a “whitening” is not easily performed, as in Sections 6.3 and 6.4, we suggest using a standard Gaussian reference distribution after centering and scaling the parameter. Future work will explore other methods for picking the reference distribution when the parameter marginal is not available.
@@ 5 Deflating score-ratio matching
Let $(\lambda_{i},\phi_{i})$, for $i=1,\dots,d$, be the eigenpairs of a symmetric matrix $A\in\mathbb{R}^{d\times d}$, with $\lambda_{1}>\lambda_{2}\geq\dots\lambda_{d}$ and $\|\phi_{i}\|_{2}=1$. Define the subunitary matrix $\Phi_{r}=[\phi_{1}\ \dots\ \phi_{r}]$ and projector $P=I-\Phi_{r}\Phi_{r}^{\top}$. Then the deflated matrix $\widetilde{A}\coloneqq PAP$ has eigenpairs $(0,\phi_{i})$ for $i=1,\dots,r$ and $(\lambda_{i},\phi_{i})$ for $i=r+1,\dots,d$.
@@
Discussion on computational costs
Here, we discuss the computational costs and scalability of the methods described in Algorithms 1 and 2. First, we emphasize that our framework functions as a pre-inference step: we only require joint prior samples, $\{x^{(j)},y^{(j)}\}_{j=1}^{N}\sim\pi_{X,Y}$. Also, rather than needing sufficient sample size and network expressivity to precisely recover the score (as one might want for posterior sampling), empirically we find that much less is needed to capture the low-dimensional structure underlying the problem, i.e., the dominant eigenspaces of the diagnostic matrices. This allows us to leverage that structure during inference, reducing the overall computational burden of the problem.
In general, approximating the score ratio encounters challenges similar to those of traditional score matching techniques. Assuming the score of the reference distribution can be evaluated efficiently, the computational cost of evaluating the traditional score matching loss and the loss function derived in Theorem 3.2 are the same up to a constant. That said, our network parameterization and regularization scheme are specifically designed to exploit low-dimensional structure in the target problem. As demonstrated in Section 6.1, our approach exhibits better sample efficiency than traditional score matching. The iterative deflating method (Algorithm 2) can further improve sample efficiency, as each score ratio function is restricted to learning a very low dimensional subspace.
@@ 6 Numerical examples
We now present several numerical experiments to show the utility of our methods. Table 1 in the Supplementary Material summarizes all network and training hyperparameter choices for each numerical example. For the problems presented in Sections 6.1 and 6.2, the score ratio is tractable and thus we can construct the true diagnostic matrices $H^{X}_{\operatorname{CDR}}$ and $H^{Y}_{\operatorname{CMI}}$. We thus evaluate the accuracy of our method by comparing the posterior approximation errors for the optimal basis transformations in (4) and (6) with the following error bounds achieved using the learned bases $\widetilde{U}$ and $\widetilde{V}$ [^3] We use a tilde to distinguish learned bases, i.e., eigenvectors of the diagnostic matrices computed using a score (ratio) network, from the eigenvectors of the exact diagnostic matrices.:
@@ 6.2 PDE-constrained inverse problem
Next, we consider an inverse problem where the forward model involves an elliptic partial differential equation (PDE). The inference parameter describes the permeability field on a two-dimensional domain $\mathcal{D}$ and the observations are pointwise measures of the corresponding pressure field. This so-called “Darcy flow” problem is a widely-used test case in the literature on nonlinear Bayesian inverse problems (Stuart, 2010; Iglesias et al., 2014; Cui et al., 2014). The permeability field $e^{\varrho}$ and the pressure field $u$ are related by the following Poisson equation (Neuman & Yakowitz, 1979; McLaughlin & Townley, 1996; Carrera et al., 2005; Sun, 2013):
@@ 6.3 Flux, a quantity-of-interest inference example
The posterior distribution of interest in this problem is $\pi_{Q|Y}$, which follows from the joint distribution $\pi_{Q,Y}$. We note in this case that evaluating the unnormalized posterior density (and thus evaluating the score function) is not tractable; more details are given in Section B.3. We seek to reduce the dimension of the observations $Y$ via Algorithm 1 using $N=10000$ samples. Figure 5(a) shows bounds on posterior approximation error that are estimated using our learned score-ratio network. The fast decay in this bound suggests that the posterior should be well approximated using relatively low-dimensional observation projections, e.g., $s=4$. To validate the utility of the corresponding learned basis $\widetilde{V}_{s}$, we perform inference using a conditional normalizing flow (specifically an unconstrained monotonic neural network (UMNN) as described in Wehenkel & Louppe (2019)) trained from samples of $\pi_{Q,Y}$; see Section B.2 for the implementation details. We build two such flows: one depending on the full-dimensional observations and the other depending on the reduced observations of dimension $s=4$ defined by the basis $\widetilde{V}_{s}$. We compare the performance of the UMNN flows learned with an increasing number of map training samples (independent from the samples used to learn the score ratio approximation) to true posterior predictive samples computed via MCMC. [^4] We perform MCMC on the high-dimensional latent parameters $X$ described in Section 6.2, and then compute log-flux samples $Q$ via (12).
@@
(a) Flux: bound on the posterior error
(b) Flux: KS statistic vs. map training set size
Figure 5: Flux problem: (a) decay of the estimated error bounds that may used to select the dimension of the reduction observation; (b) average KS statistic versus the number of training samples. Reducing the dimension of the observation enables higher-fidelity inference.
@@ 6.4 Energy price modeling
We use this dataset to estimate the score ratio for $\pi_{X|Y}$ and thus the diagnostic matrix $H^{Y}_{\operatorname{CMI}}$ via Algorithm 1. In Figure 7(a), we observe a sharp drop in the estimated posterior error from observation dimension reduction, suggesting that the leading $s=4$ modes should capture most of the information carried by the full observation about the parameter. Figures 8(b) and 8(c) plot the two leading basis vectors for the observation, and labels the observation components with the highest magnitudes. We see that the first vector primarily captures the temperature predictions at a few key cities. The second vector gives the most weight to the temporal encoding. Figure 9 plots the temperature and cloud coverage contribution of the leading four basis vectors on a map of the United States. We see that many vectors have interpretable structures. For example, the temperature component of $v_{3}$ seems to be a weighted average of the temperature predictions across the weather stations.
As before, we validate the learned reduction vectors by comparing the performance of approximate inference using full- and reduced-dimensional conditional normalizing flows as a function of the number of training samples from $\pi_{X,Y}$. While the ground-truth posterior distributions are not available in this problem, we do have the “true” realized value of the energy price. Thus we report the continuous ranked probability score (CRPS) [^5] The CRPS is defined as $\text{CRPS}(x^{*},F)=\int(F(x)-\mathds{1}_{x\geq x^{*}}){\rm d}x$ where $x^{*}$ is the true parameter value and $F$ is the cumulative distribution function of the approximate one-dimensional posterior distribution., (Gneiting & Raftery, 2007) for the approximate posterior samples, a predictive metric used to evaluate the performance of probabilistic forecasts, on average over $100$ held-out energy-price samples. Figure 7(b) shows that the reduced-dimensional flows are more predictive on average than the full-dimensional flow. We also note that the CRPS is stable with as few as $100$ training samples for the reduced-dimensional map. Figure 10 shows six histograms of posterior samples generated from full- and reduced-dimensional flows trained with $10000$ samples for different realizations of the observation variable. In two of the examples, we see that the reduced flows may capture underlying multi-modality in the posterior that the full-dimensional flow (i.e., without observation reduction) does not.
(a) Energy price: bound on the posterior error
(b) Energy price: CRPS vs. map training set size
Figure 7: Energy price problems: (a) decay of the estimated error bounds that may used to select the dimension of the reduction observation; (b) average CRPS versus the number of training samples. Reducing the dimension of the observation enables higher-fidelity inference.
@@
(a) Example posterior $1$
(b) Example posterior $2$
(c) Example posterior $3$
(d) Example posterior $4$
(e) Example posterior $5$
(f) Example posterior $6$